# Novel Application of Neutrinos to Evaluate U.S. Nuclear Weapons Performance

### v1

@@ IV.0.4 IBD Test Detector Design (FLASH)
To achieve the estimated 630 IBD neutrino event interactions in three years of pulsed running at the TAMU TRIGA reactor requires at least 20-ton fiducal mass (see calculations above). Due to mass losses in the outer veto regions and lack of efficiency, a detector with 25-tons total mass is required. The detector is to be filled with the commercially available EJ-335 scintillator oil doped with Gd [14]. This oil has proven high scintillation light output of 9600 photons/MeV of deposited energy and a thermal neutron capture time of 5.5 microseconds for a Gd concentration of 0.5% (30 microseconds for 0.1% Gd concentration). The higher the Gd concentration the shorter the capture time and better background rejection from the positron-neutron coincidence.
@@
We plan to use up to 500 standard 8” diameter PMT’s which have 22% quantum efficiency around 400 nm to detect the scintillation and Cherenkov light produced by the neutrino final state particle interactions (position and neutron capture gamma-rays). These will be instrumented with CAEN V1730 (500 MHz 14 bit charge ADCs) which will measure the neutrino interaction light pulses to nano-second time resolution and better than 5% charge resolution.
@@ IV.0.5 Detector Simulation Results
The simulation used Gd-doped EJ-335 mineral oil at various concentrations. The mineral oil filled a cylindrical detector with 200 internal PMTs facing a 7-ton fiducial volume. An incident neutrino spectrum was convoluted with the IBD capture cross section to produce an event spectrum. These events consisted of generated simultaneously with appropriate energy/momentum from a neutrino IBD event. Both initial particles were then allowed to propagate using the FTFP_BERT_HP physics list and Geant4 optical physics built from the properties of EJ-335 mineral oil scintillator [14]. Detection occurred through optical photon interaction with models of the current CCM PMTs built from LED-based single photo-electron (SPE) and radioactive source-based efficiency calibrations [27, 26].
The simulation produced expected results - see Figure 6. 100% of initial positrons were detected to some level, although not all were fully captured within the fiducial volume. The positron spectrum peaked at 3000 PEs (photo-electrons, a measure of reconstructed energy), indicating a true/reco energy ratio of 1 MeV to 1250 PEs. This gives the simulated detector sensitivity to energies between 0.01 and 30 MeV before PMT pileup becomes an issue. Approximately 80% of neutrons captured within a 10 s window at 1% Gd doping, while another simulation using 0.5% doping only captured 35% within the same 10 s window. A s thermalization time preceded the expected thermal neutron capture time of 5.5 s. The majority of these captures were on Gd-155 or Gd-157 in both cases, producing a cascade of gamma-rays totalling approximately 8 MeV. This was reconstructed in the expected energy region similar to the positron spectrum, between 8 and 10,000 PEs. The positron’s reconstructed time is taken as the neutrino arrival time since the positron is produced immediately after interaction. The neutrino energy is taken as the reconstructed positron energy plus 1.8 MeV due to interaction threshold. The energy resolution on the reconstructed positron is about 5% within the fiducial volume. This can be further reduced with advanced reconstruction techniques such as charge-time log-likelihood reconstruction analysis.
@@ IV.0.6 Additional Fundamental Science Drivers
Investigate the source of the 5 MeV bump in the reactor neutrino spectrum observed at steady state reactor experiments [29]. Measuring the time profile might yield information on the fission source responsible for the bump, though statistics in the region of interest will be a challenge.
@@
The energy and baseline of the FLASH experiment is ideal to test for sterile neutrino oscillations at region [30]. This could provide a direct measurement of neutrino oscillations consistent with Neutrino-4 and BEST results [31, 32] that have found significant evidence for or disappearance. With a planned three year run, Figure 7 shows the anticipated reach of FLASH to search for large disappearance oscillations consistent with the Neutrino-4 and BEST signals. With 2000 pulses per year, we anticipate 210 unoscillated IBD neutrinos/year after efficiency as discussed previously in Section IV.0.2. Due to the compact and movable reactor core, FLASH can perform a sensitive L/E shape analysis where many of the large systematic errors from neutrino flux and cross section only affect normalization, not shape, to first order. Also, due to the short source pulsing and IBD signal coincidence, backgrounds will be mitigated significantly and are assumed to be small - the sensitivity plot in Fig 7 includes 10% background subtraction error (see subsection IV.0.2 for details). Further details of the experiment and sensitivities will be presented in a future physics paper. The irony here is if sterile neutrinos at the do exist, they would become the largest systematic error in nuclear weapons yield extrapolation from neutrino rate measurements.
@@
A pulsed reactor produces a large number of gamma-rays and electrons that can couple to axions. Recent models have demonstrated the 1 MeV region as a motivated place to search [33]. The prompt time production of axions in the reactor core will help separate it from the IBD neutrinos coming from the slower fission decays, but will still have to contend with large prompt backgrounds from the beam neutrons and gamma-rays. The limited pulsing of the reactor will reduce signal rate, but the significant random background rejection due to the pulse timing will improve overall sensitivity. Sensitivity calculations are on-going and will be reported in a future paper.
Electromagnetic and gravitational waves currently underpin the multi-messenger era in astrophysics. These signals can disentangle the masses of stars associated with compact object collisions [34], provide insight into the nuclear equation of state [35], and indicate the production of heavy elements in explosive events [36, 37]. Neutrinos offer a third, complementary perspective on astrophysical events, giving insight into numerous processes. Because neutrinos can escape the interiors of astrophysical environments, they allow observation of internal mechanisms and conditions [38]. Precision energy- and time-resolved measurements of pulsed neutrinos in terrestrial environments, as suggested in this paper, will help to refine the prediction of neutrino signals for astrophysics. A well characterized neutrino signal can enhance understanding of astrophysical nuclear processes. Neutrino measurements can provide constraints on standard model physics and extensions thereto, which directly affect the evolution of extreme astrophysical environments.
@@ V Conclusions
Due to neutrinos prolific production in fission reactions during a nuclear weapons test, they can be used as a novel and complementary tool for diagnostics. With large fiducial mass and prolific scintillation light detection, the modern neutrino detector using the IBD reaction can easily detector thousands of neutrinos from a stand off distance of 500 meters. Should the U.S. every go back to nuclear weapons testing, the use of neutrinos would provide a new tool to better understand and diagnose the performance of a weapon. In the mean time, the technology and methods used can be tested at a pulsed reactor such as the TAMU TRIGA reactor, which is a good surrogate for a nuclear weapons test. This enables studying yield extraction, non-equilibrium fission rates, and systematic errors. As well, tests can be performed of the neutrino detector technology in a pulsed environment, such as neutrino pileup, readout rates, triggering, etc. Finally, with a modest sized 20-ton fiducial mass scintillation detector (FLASH) deployed at a pulsed reactor, tests of the Neutrino-4 and BEST disappearance signal can be performed, which could be indication of sterile neutrino oscillations at the scale. Using a reactor in pulsed mode for neutrino physics is novel and would yield significant background reduction and enable the study of the time evolution of fission neutrinos. Such a rich program is of modest cost and could be deployed in a few years.

### v2

@@ IV.0.4 IBD Test Detector Design (FLASH)
To achieve the estimated 654 IBD antineutrino event interactions in three years of pulsed running at the TAMU TRIGA reactor requires at least a 20-ton fiducal mass (see calculations above). Due to mass losses in the outer veto regions and lack of efficiency, a detector with 25-tons total mass is required. The detector is to be filled with the commercially available EJ-335 scintillator oil doped with Gd [29]. This oil has proven high scintillation light output of 9600 photons/MeV of deposited energy and a thermal neutron capture time of 5.5 microseconds for a Gd concentration of 0.5% (30 microseconds for 0.1% Gd concentration). The higher the Gd concentration the shorter the capture time and better background rejection from the positron-neutron coincidence.
@@
We plan to use up to 500 standard 8” diameter PMT’s which have 22% quantum efficiency around 400 nm to detect the scintillation light produced by the IBD final state particle interactions (position and neutron capture gamma-rays). These will be instrumented with CAEN V1730 (500 MHz 14 bit charge ADCs) which will measure the antineutrino interaction light pulses to nanosecond time resolution and better than 5% charge resolution.
@@ IV.0.5 Detector Simulation Results
The simulation used Gd-doped EJ-335 mineral oil at various concentrations. The mineral oil filled a cylindrical detector with 200 internal PMTs facing a 7-ton fiducial volume. An incident antineutrino spectrum was convoluted with the IBD capture cross section to produce an event spectrum. These events consisted of generated simultaneously with appropriate energy/momentum from a antineutrino IBD event. Both initial particles were then allowed to propagate using the FTFP_BERT_HP physics list and Geant4 optical physics built from the properties of EJ-335 mineral oil scintillator [29]. Detection occurred through optical photon interaction with models of the current CCM PMTs built from LED-based single photo-electron (SPE) and radioactive source-based efficiency calibrations [32, 31].
The simulation produced expected results - see Figure 6. 100% of initial positrons were detected to some level, although a small fraction partially escaped the fiducial volume adding to the energy smearing. The positron spectrum peaked at 3000 PEs (photo-electrons, a measure of reconstructed energy), indicating a true/reco energy ratio of 1 MeV to 1250 PEs. This gives the simulated detector sensitivity to energies between 0.01 and 30 MeV before PMT non-linearity and saturation may become an issue. We expected a s neutron thermalization time and a s thermal neutron capture time for 1% Gd doping from similar processes in water [34, 35], with a longer capture time for lower concentrations of Gd. The simulation demonstrated this behavior. Approximately 80% of neutrons captured within a 10 s window at 1% Gd doping, while using 0.5% doping only captured 35% within the same window. The majority of captures were on Gd-155 or Gd-157 in both cases, producing a cascade of gamma-rays totaling approximately 8 MeV. This was reconstructed in the expected energy region compared to the positron spectrum, between 8,000 and 10,000 PEs, with the spread coming from 20% energy resolution smearing. This high uncertainty is derived from the multi-particle cascade, sometimes producing much higher reconstructed energy up to 35000 PE 30 MeV. A fraction of the captures were on H, producing a second feature in Figure 6 at MeV and 6 s. The positron’s reconstructed time is taken as the antineutrino arrival time since the positron is produced immediately after interaction. The antineutrino energy is taken as the reconstructed positron energy plus 1.8 MeV due to interaction threshold. The energy resolution on the reconstructed positron is about 5% within the fiducial volume. This can be further reduced with advanced reconstruction techniques such as charge-time log-likelihood reconstruction analysis. For the antineutrino rate calculations in this paper we assume 1% Gd doping giving an overall IBD detection efficiency of 80%.
@@ IV.0.6 Additional Fundamental Science Drivers
Investigate the source of the 5 MeV bump in the reactor antineutrino spectrum observed at steady state reactor experiments [36]. One interesting measurement is the time profile might yield information on a prompt new BSM process that can be separated in time from slow fission antineutrino sources. However, statistics will be a challenge for FLASH since the bump is only 10% of the total neutrino rate and is more a venue for future research.
@@
The energy and baseline of the FLASH experiment is ideal to test for sterile neutrino oscillations at m/MeV region [37]. This could provide a direct measurement of neutrino oscillations consistent with Neutrino-4 and BEST results [38, 39] that can be interpreted as or disappearance. With a planned three year run, Figure 7 shows a toy model anticipated reach of FLASH to search for large disappearance oscillations consistent with the Neutrino-4 and BEST signals. With 2000 pulses per year and detector at 4 m, we anticipate 218 unoscillated IBD antineutrinos/year after efficiency as discussed previously in Section IV.0.2. Due to the compact core, FLASH can perform a sensitive L/E shape analysis where many of the large systematic errors from antineutrino flux and cross section only affect normalization, not shape, to first order. Also, due to the short source pulsing and IBD signal coincidence, backgrounds will be mitigated significantly and are assumed to be small - the sensitivity plot in Fig 7 includes 10% background subtraction error (see subsection IV.0.2 for details). Other errors included is the position smearing effects from source size of 0.5 m on each side (standard deviation of 0.15 m) and event position reconstruction of 10 cm position resolution. Finally, the IBD spectrum energy is smeared by 5% to account for the detector energy reconstruction resolution. If an oscillation signal is observed, then the movable reactor core position could be changed to optimize sensitivity at the best fit point. Further details of the experiment and sensitivities will be presented in a future physics paper. The irony here is if sterile neutrinos at the do exist, they would become the largest systematic error in nuclear weapons yield extrapolation from neutrino rate measurements.
@@
A pulsed reactor, and nuclear weapon blast, produces a large number of prompt (less than a few hundred nanoseconds) gamma-rays and electrons [40] that can potentially couple to axions or axion like particles (ALP). Recent models have demonstrated the 1 MeV region as a motivated place to search [41]. The prompt time production of axions/ALPs in the reactor core will help separate it from the IBD antineutrinos coming from the slower fission decays, but will still have to contend with large prompt backgrounds from the beam neutrons and gamma-rays. The limited pulsing of the reactor will reduce signal rate, but the significant random background rejection due to the pulse timing will improve overall sensitivity. Sensitivity calculations are on-going and will be reported in a future paper.
@@ V Conclusions
Due to the prolific production of neutrinos in fission reactions during a nuclear weapons test, they can be used as a novel and complementary tool for diagnostics. With large fiducial mass and prolific scintillation light detection, modern neutrino detector using the IBD reaction can easily detect thousands of antineutrinos from a stand off distance of 500 meters. Should the U.S. ever go back to nuclear weapons testing, the use of neutrinos would provide a new tool to better understand and diagnose the performance of a weapon. In the meantime, the technology and methods used can be tested at a pulsed reactor such as the TAMU TRIGA reactor, which is a good surrogate for a nuclear weapons test. This enables studying yield extraction, non-equilibrium fission rates, and systematic errors. As well, tests can be performed of the IBD neutrino detector performance in a pulsed environment, such as neutrino pileup, readout rates, triggering, reconstruction, etc. Finally, with a modest sized 20-ton fiducial mass scintillation detector (FLASH) deployed at a pulsed reactor, tests of the Neutrino-4 and BEST disappearance signal can be performed, which could be indication of sterile neutrino oscillations at the scale. Using a reactor in pulsed mode for neutrino physics is novel and would yield significant background reduction and enable the study of the time evolution of fission neutrinos. Such a rich program is of modest cost and could be deployed in a few years.